OpenStudy (anonymous):

At time t in seconds, a particle's distance s(t), in centimeters, from a point is given by s(t) = 4 + 3sin(t) What is the average velocity of the particle from t = π/3 to t = 7π/3?

OpenStudy (anonymous):

The average velocity of a position function is the same as "the average rate of change of the position function." Over an interval \([a,b]\), the average rate of change of a function \(s(t)\) is given by \[\frac{s(b)-s(a)}{b-a}\]

OpenStudy (anonymous):

So in this case, the avg rate of change is \[\frac{(4+\sin3\left(\frac{7\pi}{3}\right))-(4+\sin3\left(\frac{\pi}{3}\right))}{\frac{7\pi}{3}-\frac{\pi}{3}}\]

OpenStudy (anonymous):

thanks! can you help me simplify that?

OpenStudy (campbell_st):

I'll do the denominator its \[2 \pi\]

OpenStudy (anonymous):

\(\sin3\left(\dfrac{7\pi}{3}\right)=\sin7\pi=0\). Do the same with the other trig term.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw